OpenStudy (anonymous):

Simplify. 3 times the third root of 24

OpenStudy (anonymous):

\[3 \sqrt[3]{24}\] That's how it looks written out

OpenStudy (anonymous):

8.65

OpenStudy (anonymous):

These are my options: 6 times the third root of 3 18 9 times the third root of 3 27

OpenStudy (mertsj):

\[3\sqrt[3]{24}=3\sqrt[3]{8(3)}=3\sqrt[3]{8}\sqrt[3]{3}=3(2)\sqrt[3]{3}=6\sqrt[3]{3}\]

OpenStudy (anonymous):

Perfect! Thanks!

OpenStudy (mertsj):

yw